### v1

# On real forms of Fermat hypersurfaces

Abstract.
In this paper, we compute the number of real forms of Fermat hypersurfaces for degree $d\geq 3$ except the degree $4$ surface case, and give explicit descriptions of them.
1. Introduction
A real form of a complex algebraic variety $X$ is a real algebraic variety whose complexification is isomorphic to $X$. Now, some examples of smooth complex varieties with infinitely many nonisomorphic real forms are known. The first example was obtained by Lesieutre in [Les18] for varieties of dimension $6$, and the existence of such varieties for any dimension $n\geq 2$ was shown in [DO19]. Also, for dimension $1$, it is known that there are no such varieties. In addition, the construction of smooth complex rational affine surface with uncountably many nonisomorphic real forms, and some other results about finiteness problem of real forms is summarized in [Bot21].
Considering the real forms of specific varieties is also an interesting question. Fermat hypersurface of dimension $n$ and degree $d$ is a projective variety defined as follows:
$F^{n}_{d(,k)}:=\left(X_{0}^{d}+\cdots+X_{n}^{d}+X_{n+1}^{d}=0\right)\subset\mathbb{P}_{k}^{n+1}$
where $k$ is the base field, and we don’t write it explicitly when it is clear. In [Ogu], the following question is asked:
Question 1.1.
Find all real forms of $F^{n}_{d,\mathbb{C}}$ up to $\mathbb{R}$-isomorphisms.
When $d=1$ or $(n,d)=(1,2)$, $F^{n}_{d}$ is isomorphic to a projective space, and this has exactly $1$ real form if $n$ is even, and $2$ real forms if $n$ is odd, up to isomorphisms, which is proved in [Ben16]. Also, it is proved in [Rus02] that $F^{2}_{2}$ has exactly $4$ real forms. Note that by [BS64], for degree $2$, $F^{n}_{2}$ has only finitely many real forms up to isomorphisms as the first Galois cohomology of an affine linear algebraic group is finite.
In this paper, we give the following partial answer to the above question:
Main Theorem.
Suppose $(n,d)\neq(2,4)$ and $d\geq 3$. Then, the number of real forms of $F^{n}_{d,\mathbb{C}}$ is
$\begin{cases}\left\lfloor\frac{n}{2}\right\rfloor+2&(d\equiv 1\pmod{2})\\ \frac{(n+3)(n+5)}{8}&(d\equiv 0,\ n\equiv 1\pmod{2})\\ \frac{(n+4)(n+6)}{8}&(d\equiv 0,\ n\equiv 0\pmod{2})\end{cases}$
We prove this theorem by computing $H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})$. In Section 3, we deal with the case $(n,d)\neq(1,3),(2,4)$, $d\geq 3$, and in Section 4, we deal with the case $(n,d)=(1,3)$. In both cases, we have already known the automorphism group $\operatorname{\operatorname{Aut}}F^{n}_{d}$, as the former case is due to Shioda [Shi87], and the latter is due to the fact that $F^{1}_{3}$ is elliptic. So we will check the condition for an automorphism of $F^{n}_{d}$ to be a cocycle, compute a conjugation of a cocycle by an automorphism, and get this result.
We also give explicit descriptions of these real forms for the curve case in Theroem 5.1, and for the surface case in Theorem 5.2 in the way we can generalize it to higher dimensions, and give a topological description of a real locus of some real forms (in Theorem 5.3, and in particular, Theorem 5.4 for the curve case). We also show that $F_{4}^{2}$ has at least $4$ non-isomorphic real forms in Remark 5.5.
2. Preliminary about real forms
First, we review the definition of a real form and related concepts.
Definition 2.1.
Let $X$ be a complex algebraic variety and $X^{\prime}$ a real algebraic variety.
  1. (1)
    The complexification of $X^{\prime}$ is the complex algebraic variety
    ${X^{\prime}}_{\mathbb{C}}:=X^{\prime}\times_{\operatorname{Spec}\mathbb{R}}\operatorname{Spec}\mathbb{C}.$
  2. (2)
    A real form of $X$ is a real algebraic variety $X_{\mathbb{R}}$ with a $\mathbb{C}$-isomorphism $(X_{\mathbb{R}})_{\mathbb{C}}\xrightarrow[]{\sim}X$.
  3. (3)
    A real structure on $X$ is an anti-involution $\rho\colon X\to X$, i.e., an involution that makes the following diagram commute:
[…]
    where the bottom arrow is the morphism induced by the complex conjugation.
  4. (4)
    Real structures $\rho$, $\rho^{\prime}$ on $X$ are equivalent if there exists a $\varphi\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ such that $\rho^{\prime}=\varphi^{-1}\circ\rho\circ\varphi$.
Theorem 2.2 ([BS64]).
Let $X$ be a complex quasi-projective variety with a real structure $\rho$ on $X$. Then, there are bijections between
  • the set of $\mathbb{R}$-isomorphism classes of real forms of $X$,
  • the set of equivalent classes of real structures on $X$, and
  • $H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}_{\mathbb{C}}X)$ where the nontrivial element of $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ acts on $\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ by the conjugation of $\rho$.
It is clear by definition that there exists a bijection between the second and the third set, and existence of a bijection between the first and the third set is proved in [Se02, Chapter III, 1.3].
Remark 2.3.
In the above condition,
  • $\alpha\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ is a cocycle if and only if $(\alpha\circ\rho)^{2}=\operatorname{id}_{X}$.
  • cocycles $\alpha,\beta\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ are equivalent if and only if there exists $\varphi\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ such that $\beta\circ\rho=\varphi^{-1}\circ(\alpha\circ\rho)\circ\varphi$, and we denote it by $\alpha\sim\beta$.
3. The number of real forms of Fermat hypersurfaces
Before thinking about real forms of Fermat hypersurfaces, we want to think about automorphisms of them. Clearly, there are two types of automorphisms when the base field is algebraically closed and of characteristic 00. An automorphism of the first type is the one with which we can identify $\sigma\in S_{n+2}$ defined as follows:
$F^{n}_{d}\to F^{n}_{d};[X_{0}:\cdots:X_{n+1}]\mapsto[X_{\sigma(0)}:\cdots:X_{\sigma(n+1)}]$
with $\sigma$ viewed as a permutation of $\{0,\ldots,n+1\}$.
An automorphism of the second type is the one denoted by $g(i,j)$ for $i,j\in\mathbb{Z}$, $10\leq i\leq n+$ 1 and defined as follows:
$F^{n}_{d}\to F^{n}_{d};[X_{0}:\cdots X_{n+1}]\mapsto[X_{0}:\cdots:\xi^{j}X_{i}:\cdots:X_{n+1}]$
with $\xi=\xi_{d}$ a primitive $d$-th root of unity.
Let $G^{n}_{d}$ be the subgroup of $\operatorname{\operatorname{Aut}}F^{n}_{d}$ generated by $g(i,j)$ for all $i,j\in\mathbb{Z}$. Then, the following result is known.
Theorem 3.1 ([Shi87]).
Let $k$ be an algebraically closed field with $\operatorname{char}k=0$, and suppose $(n,d)\neq(1,3),(2,4)$ and $d\geq 3$. Then,
$\operatorname{\operatorname{Aut}}F^{n}_{d,k}=G^{n}_{d}\cdot S_{n+2}$
where normal-$\cdot$ denotes a semidirect product.
In particular, $\operatorname{\operatorname{Aut}}F^{n}_{d,\mathbb{C}}=G^{n}_{d}\cdot S_{n+2}$.
Though this theorem is proved only in the curve case in that paper, the same argument works in higher dimensions as we can show that every automorphism is linear.
From now on, we assume $(n,d)\neq(1,3),(2,4)$, $d\geq 3$, and the base field is $\mathbb{C}$ unless otherwise stated.
Let $\tilde{\rho}\colon\mathbb{P}^{n+1}\to\mathbb{P}^{n+1}$ be the complex conjugation, i.e., the $\mathbb{R}$-morphism defined by $\mathbb{C}[X]\to\mathbb{C}[X]$ that sends $X_{i}$ to $X_{i}$ for all $i$ and $c\in\mathbb{C}$ to $\bar{c}$. This is a real structure of $\mathbb{P}^{n+1}$, and it induces the real structure $\rho\colon F^{n}_{d}\to F^{n}_{d}$ on $F^{n}_{d}$.
Remark 3.2.
By straightforward calculations, for $i,j\in\mathbb{Z}$, $10\leq i\leq n+$ 1, $\sigma\in S_{n+2}$,
  • $g(i,j)\circ\sigma=\sigma\circ g(\sigma^{-1}(i),j)$.
  • $\sigma\circ\rho=\rho\circ\sigma$.
  • $g(i,j)\circ\rho=\rho\circ g(i,d-j)=\rho\circ g(i,j)^{-1}$.
The following result is the main step to prove the main theorem.
Proposition 3.3.
  1. (1)
    $\alpha=\mu\sigma\in\operatorname{\operatorname{Aut}}F^{n}_{d}$, with $\mu=g(0,p_{0})\cdots g(n+1,p_{n+1})\in G_{d}^{n}$, $\sigma\in S_{n+2}$, is a cocycle if and only if the following two conditions hold:
    • $\sigma$ is of type $1^{l}\cdot 2^{m}$ with $l+2m=n+2$.
    • $p_{i}\equiv p_{\sigma(i)}\pmod{d}$ if $i\in\operatorname{supp}\sigma$.
    Here, $\operatorname{supp}\sigma$ is the support of a permutation $\sigma$, i.e., $\operatorname{supp}\sigma:=\left\{i|\sigma(i)\neq i\right\}$.
  2. (2)
    Cocycles $\beta=\lambda\tau$, $\gamma$ with $\lambda\in G^{n}_{d}$ and $\tau\in S_{n+2}$ are equivalent if and only if there exist $\psi\in S_{n+2}$ and $q_{j}\in\mathbb{Z}$ for $10\leq j\leq n+$ 1 with $q_{j}\equiv q_{\sigma(j)}\pmod{2}$ for all $j$ and $q_{l}\equiv 0\pmod{2}$ for all $l\notin\operatorname{supp}\tau$ such that
    $\gamma=\psi^{-1}\cdot\left(g(0,q_{0})\cdots g(n+1,q_{n+1})\beta\right)\cdot\psi.$
Proof.
(1) By definition, $\alpha$ is a cocycle if and only if $(\mu\sigma\rho)^{2}=\operatorname{id}$. Also,
$(\mu\sigma\rho)^{2}$ $=\mu\sigma\mu^{-1}\sigma$
$=g(0,{p^{\prime}}_{0})\cdots g(n+1,{p^{\prime}}_{n+1})\cdot\sigma^{2}$
where
${p^{\prime}}_{i}=\begin{cases}0&(i\notin\operatorname{supp}\sigma)\\ p_{i}-p_{\sigma(i)}&(i\in\operatorname{supp}\sigma)\end{cases}$
Therefore, $(\mu\sigma\rho)^{2}=\operatorname{id}$ is equivalent to $\sigma^{2}=\operatorname{id}$ and ${p^{\prime}}_{i}\equiv 0\pmod{d}$, and the result follows.
(2) Suppose $\beta=\lambda\tau$ be a cocycle and $\varphi=\nu\psi\in\operatorname{\operatorname{Aut}}F^{n}_{d}$ with $\lambda,\nu=g(0,r_{0})\cdots g(n+1,r_{n+1})\in G^{n}_{d},\ \tau,\psi\in S_{n+2}$. Then,
$\varphi^{-1}\circ\beta\rho\circ\varphi$
$=$ $(\psi^{-1}\nu^{-1})\cdot(\lambda\tau\rho)\cdot(\nu\psi)$
$=$ $\psi^{-1}\cdot(\nu^{-1}\lambda\tau\nu^{-1})\cdot\psi\rho$
$=$ $\psi^{-1}\cdot(\nu^{-1}\lambda\cdot g(\tau(0),d-r_{0})\cdots g(\tau(n+1),d-r_{n+1}))\cdot\tau\psi\rho$
$=$ $\psi\cdot(g(0,s_{0})\cdots g(n+1,s_{n+1})\cdot\lambda\tau)\cdot\psi\rho$
where
$s_{i}=\begin{cases}-2r_{i}&(i\notin\operatorname{supp}\tau)\\ -r_{i}-r_{\tau(i)}&(i\in\operatorname{supp}\tau)\end{cases}$
As $r_{i}$ varies for all integers, the result follows. ∎
Note that, as cocycles,
$I_{r}=\xi^{m}I_{r}$
for $m\in\mathbb{Z}$ by definition, and
$\begin{pmatrix}1&\\ &I_{r}\end{pmatrix}\sim\begin{pmatrix}\xi^{2i}&\\ &I_{r}\end{pmatrix},\ \begin{pmatrix}I_{r}&&\\ &0&1\\ &1&0\end{pmatrix}\sim\begin{pmatrix}I_{r}&&\\ &0&\xi^{j}\\ &\xi^{j}&0\end{pmatrix}$
for $i,j\in\mathbb{Z}$ by the proposition.
By this proposition, we can describe $H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})$.
Theorem 3.4.
Suppose $d$ is odd. Then, any cocycles are equivalent to exactly one of the following forms:
[…]
In particular, $F^{n}_{d}$ has exactly $\left\lfloor\frac{n}{2}\right\rfloor+2$ real forms.
In particular, if $d\geq 3$ is odd, then $F^{1}_{d}$ ($d\neq 3$) has exactly $2$ real forms and $F^{2}_{d}$ has $3$ real forms.
Proof.
As $d$ is odd,
$\begin{pmatrix}1&\\ &I_{r}\end{pmatrix}\sim\begin{pmatrix}\xi^{i}&\\ &I_{r}\end{pmatrix}$
for all $i$, and the former statement follows from Proposition 3.3. Also, as any two of $H^{r}_{n}$’s are not equivalent,
$H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})=\left\{H^{r}_{n}|0\leq r\leq n+2\right\}$
where we identify cocycles with their equivalent classes, and the latter statement follows. ∎
Theorem 3.5.
Suppose $d$ is even. Then, any cocycles are equivalent to exactly one of the forms:
$K^{s,t}_{n}:=\begin{pmatrix}\xi I_{s}&&&&&&\\ &I_{t}&&&&&\\ &&0&1&&&\\ &&1&0&&&\\ &&&&\ddots&&\\ &&&&&0&1\\ &&&&&1&0\end{pmatrix}\ (s,t\in\mathbb{Z}_{\geq 0},\ s+t\leq n+2)$
Also, $F^{n}_{d}$ has exactly $\frac{(n+3)(n+5)}{8}$ real forms for $n$ odd, and $\frac{(n+4)(n+6)}{8}$ real forms for $n$ even.
In particular, if $d\geq 3$ is even, then $F^{1}_{d}$ has exactly $3$ real forms and $F^{2}_{d}$ ($d\not=4$) has $6$ real forms.
Proof.
The former statement is clear from proposition 3.3.
By the proposition, for $s,s^{\prime},t,t^{\prime}\in\mathbb{Z}$ with $s+t=s^{\prime}+t^{\prime}$, $K^{s,t}_{n}\sim K^{s^{\prime},t^{\prime}}$ if and only if there exists a permutation matrix $\sigma$ and $a_{1}$, $\ldots$, $a_{n+2}$ with $a_{1}\equiv\cdots\equiv a_{m}\pmod{2}$ such that
$K^{s,t}_{n}\cdot\left(\sigma^{-1}K^{s^{\prime},t^{\prime}}_{n}\sigma\right)^{-1}=\begin{pmatrix}\xi^{a_{1}}&&\\ &\ddots&\\ &&\xi^{a_{n+2}}\end{pmatrix}.$
This condition is equivalent to $s=s^{\prime}$ or $s=t^{\prime}$.
Therefore,
[…]
where we identify cocycles with their equivalent classes. From this, when $n=2k-1$ for $k\in\mathbb{Z}$, the number of real forms is
$\sum_{i=0}^{k}(i+1)=\frac{(k+1)(k+2)}{2}=\frac{(n+3)(n+5)}{8},$
and when $n=2k$ for $k\in\mathbb{Z}$, the number of real forms is
$\sum_{i=0}^{k+1}(i+1)=\frac{(k+2)(k+3)}{2}=\frac{(n+4)(n+6)}{8},$
so the result follows. ∎
We omit subindexes of $H^{r}_{n}$ or $K^{s,t}_{n}$ if it is clear from the context.
4. The elliptic curve case
In this section, we consider the real forms of $F^{1}_{3}=(X_{0}^{3}+X_{1}^{3}+X_{2}^{3}=0)\subset\mathbb{P}^{2}$. First, the substitution
$X=\frac{-12X_{2}}{X_{0}+X_{1}},\ Y=\frac{36(X_{0}-X_{1})}{X_{0}+X_{1}}$
gives an elliptic curve $Y^{2}=X^{3}-432$ with $j$-invariant 00, and this can be identified with $E:=\mathbb{C}/(\mathbb{Z}+\zeta\mathbb{Z})$ where $\zeta$ is a primitive $6$-th root of unity, which we fix in this section.
Then, any automorphisms of $E$ can be expressed by
$z\mapsto\zeta^{i}z+a$
with $i\in\mathbb{Z}$ and $a$ in the fundamental domain (cf. [Sil09]). Note that the real structure $\rho\colon F^{1}_{3}\to F^{1}_{3}$ corresponds to $E\to E$ that maps $z$ to $\bar{z}$ where we identify $z\in\mathbb{C}$ with the image of $z$ by the canonical surjection $\mathbb{C}\to\mathbb{C}/(\mathbb{Z}+\zeta\mathbb{Z})$. We also denote it by $\rho\colon E\to E$.
Proposition 4.1.
We fix $\xi\in\mathbb{C}$ so that $\xi^{2}=\zeta$.
  1. (1)
    $\alpha\colon E\to E;\ z\mapsto\zeta^{i}z+a$ is a cocycle if and only if there exists $r\in\mathbb{R}$ such that $a=\sqrt{-1}\xi^{i}r$.
  2. (2)
    Any cocycles are equivalent to $z\mapsto z$ or $z\mapsto-z$, which are not equivalent to each other.
Proof.
(1) By definition, $\alpha$ is a cocycle if and only if $(\alpha\circ\rho)^{2}=\operatorname{id}$.
$(\alpha\circ\rho)^{2}(z)$ $=\zeta^{i}(\overline{\zeta^{i}\bar{z}+a})+a$
$=z+(\zeta^{i}\bar{a}+a)$
Therefore, the result follows.
(2) Let $\sigma\colon E\to E;\ z\mapsto\zeta^{i}z+a$ be a cocycle and $\tau\colon E\to E;\ z\mapsto\zeta^{j}z+b$. Then, $\tau^{-1}\colon z\mapsto\zeta^{-j}z-\zeta^{j}b$. Therefore,
$(\tau^{-1}\circ(\sigma\rho)\circ\tau)(z)$
$=$ $\tau^{-1}(\zeta^{i}(\overline{\zeta^{j}z+b})+a)$
$=$ $\zeta^{i-2j}\bar{z}+\zeta^{-j}a-\zeta^{j}b+\zeta^{i-j}\bar{b}$
First, $z\mapsto z$ and $z\mapsto-z$ are not equivalent to each other by this calculation.
Also, in the above, we can write $a=\sqrt{-1}\xi^{i}x$ for $x\in\mathbb{R}$ by (1). Then, by setting
$j=0,\ b=\frac{1+\sqrt{-1}}{2}\xi^{i}x,$
we see that $z\to\zeta^{i}z+a$ is equivalent to $z\to\zeta^{i}z$, so the result follows. ∎
Note that the negation map in the above corresponds to a linear automorphism
$F^{1}_{3}\to F^{1}_{3};\ [X_{0}:X_{1}:X_{2}]\mapsto[X_{1}:X_{0}:X_{2}]$
by taking $[1:-1:0]$ as a unit. From this, we immediately get the following result:
Theorem 4.2.
In $F^{1}_{3}$, any cocycles are equivalent to $H^{3}_{1}$ or $H^{1}_{1}$.
In particular, $F^{1}_{3}$ has exactly $2$ real forms.
5. Explicit description of real forms of Fermat hypersurfaces
First, we consider the curve case $F^{1}_{d}=(X_{0}^{d}+X_{1}^{d}+X_{2}^{d}=0)$.
Theorem 5.1.
Assume that $d\geq 3$. Then, the correspondence between cocycles and real forms of $F^{1}_{d}$ is given by the following table.
cocycle real form
$K^{0,3}_{1}$ (or $H^{0}_{1}$) $(Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}=0)\subset\mathbb{P}^{2}_{\mathbb{R}}$
$K^{1,2}_{1}$ $(-Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}=0)\subset\mathbb{P}^{2}_{\mathbb{R}}$
$K^{0,1}_{1}$ (or $H^{2}_{1}$) $(Y_{0}^{d}+(Y_{1}+\sqrt{-1}Y_{2})^{d}+(Y_{1}-\sqrt{-1}Y_{2})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{2}$
Proof.
First, the real form corresponding to $K^{0,3}$ is clear.
Next, we consider the real form corresponding to $K^{1,2}$. The invariant ring of the homomorphism $\alpha\colon\mathbb{C}[X_{0},X_{1},X_{2}]\to\mathbb{C}[X_{0},X_{1},X_{2}]$ that sends $c\in\mathbb{C}$ to $\bar{c}$, $X_{0}$ to $\xi_{d}X_{0}$, $X_{1}$ to $X_{1}$, and $X_{2}$ to $X_{2}$ is $\mathbb{R}[\xi_{2d}X_{0},X_{1},X_{2}]$ with $\xi_{2d}^{2}=\xi_{d}$.
In fact, for all $f(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$, there exists a unique polynomial $\mu(f)(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$ such that
$f(X_{0},X_{1},X_{2})=\mu(f)(\xi_{2d}X_{0},X_{1},X_{2}),$
and $\alpha(f)=f$ if and only if $\mu(f)$ is a real coefficient polynomial, as $\xi_{2d}X_{0}$, $X_{1}$, $X_{2}$ are invariant under $\alpha$.
Therefore, by setting
$Y_{0}:=\xi_{2d}X_{0},\ Y_{1}:=X_{1},\ Y_{2}:=X_{2},$
we see that the corresponding real form is
$(-Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{2}.$
Finally, we consider the real form corresponding to $K^{1,2}$. The invariant ring of the homomorphism $\beta\colon\mathbb{C}[X_{0},X_{1},X_{2}]\to\mathbb{C}[X_{0},X_{1},X_{2}]$ that sends $c\in\mathbb{C}$ to $\bar{c}$, $X_{0}$ to $X_{0}$, $X_{1}$ to $X_{2}$ and $X_{2}$ to $X_{2}$ is $\mathbb{R}[X_{0},X_{1}+X_{2},(1-\sqrt{-1})(X_{1}-\sqrt{-1}X_{2})]$.
We can show this by a similar argument as above. In fact, for all $f(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$, there exists a unique polynomial $\nu(f)(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$ such that
$f(X_{0},X_{1},X_{2})=\nu(f)(X_{0},X_{1}+X_{2},(1-\sqrt{-1})(X_{1}-\sqrt{-1}X_{2})),$
and $\beta(f)=f$ if and only if $\nu(f)$ is a real coefficient polynomial.
Therefore, by setting
$Y_{0}:=X_{0},\ {Y^{\prime}}_{1}:=X_{1}+X_{2},\ {Y^{\prime}}_{2}:=(1-\sqrt{-1})(X_{1}-\sqrt{-1}X_{2}),$
and also, setting
$Y_{1}:=\frac{{Y^{\prime}}_{1}}{2},\ Y_{2}:=\frac{{Y^{\prime}}_{2}-{Y^{\prime}}_{1}}{2},$
we see that the corresponding real form is
$(Y_{0}^{d}+(Y_{1}+\sqrt{-1})^{d}+(Y_{1}-\sqrt{-1})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{2}.$
In the surface case, by using the same argument as above, we get the following result:
Theorem 5.2.
Assume $d\geq 3,\neq 4$. Then, the correspondence between cocycles and real forms of $F^{2}_{d}$ is given by the following table.
cocycle real form
$K^{0,4}_{2}$ (or $H^{4}_{2}$) $(Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
$K^{1,3}_{2}$ $(-Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
$K^{2,2}_{2}$ $(-Y_{0}^{d}-Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
$K^{0,2}_{2}$ (or $H^{2}_{2}$) $(Y_{0}^{d}+Y_{1}^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
$K^{1,1}_{2}$ $(-Y_{0}^{d}+Y_{1}^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
$K^{0,0}_{2}$ (or $H^{0}_{4}$) $((Y_{0}+\sqrt{-1}Y_{1})^{d}+(Y_{0}-\sqrt{-1}Y_{1})^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
In the same way, we can extend it to the general case. In particular, all real forms of $F_{d}^{n}$ are the $n$-dimensional hypersurface of degree $d$.
Next, we will give a topological description of a real locus of these real forms.
Theorem 5.3.
  1. (1)
    The real locus of
    $(X_{0}^{d}+\cdots+X_{n+1}^{d}=0)$
    with $d$ even is $\emptyset$.
  2. (2)
    The real locus of
    $(X_{0}^{d}+\cdots+X_{n}^{d}-X_{n+1}^{d}=0)$
    with $d$ even is homeomorphic to $S^{n}$.
  3. (3)
    The real locus of
    $(X_{0}^{d}+\cdots+X_{i}^{d}-X_{i+1}^{d}-\cdots-X_{n+1}^{d}=0)$
    with $d$ even is homeomorphic to $\mathbb{P}^{i}(\mathbb{R})\times\mathbb{P}^{n-i}(\mathbb{R})$.
  4. (4)
    The real locus of
    $\left(\sum_{k=0}^{i-1}X_{k}^{d}+\sum_{l=0}^{\frac{n-i}{2}}\left(\left(X_{i+2l}+\sqrt{-1}X_{i+2l+1}\right)^{d}+\left(X_{i+2l}-\sqrt{-1}X_{i+2l+1}\right)^{d}\right)=0\right)$
    with $d$ odd and $i>0$ is homeomorphic to $\mathbb{P}^{n}(\mathbb{R})$.
  5. (5)
    The real locus of
    $(X_{0}^{d}+\cdots+X_{n-1}^{d}+(X_{n}+\sqrt{-1}X_{n+1})^{d}+(X_{n}-\sqrt{-1}X_{n+1})^{d}=0)$
    with $d$ even is homeomorphic to $\underbrace{S^{n}\sqcup\cdots\sqcup S^{n}}_{\frac{d}{2}}$.
Proof.
In each case, let $Z$ be the real locus.
(1) is clear.
(2) Note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $x_{n+1}\neq 0$. Therefore, we can define a continuous map $Z\to S^{n}$ as follows:
$[x_{0}:\cdots:x_{n}:1]\mapsto\frac{1}{\sqrt{x_{0}^{2}+\cdots+x_{n}^{2}}}(x_{0},\ldots,x_{n}).$
As it is a bijection, it is also a homeomorphism, and thus, $Z\approx S^{n}$.
(3) Note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{0},\ldots,x_{i})\neq(0,\ldots,0)$, $(x_{i+1},\ldots,x_{n+1})\neq(0,\ldots,0)$. Thus, we can define a continuous map $Z\to\mathbb{P}^{i}(\mathbb{R})\times\mathbb{P}^{n-i}(\mathbb{R})$ as follows:
$[x_{0}:\cdots:x_{n+1}]\mapsto([x_{0}:\cdots:x_{i}],[x_{i+1}:\cdots:x_{n+1}]).$
As it is a bijection, it is also a homeomorphism, and thus, $Z\approx\mathbb{P}^{i}(\mathbb{R})\times\mathbb{P}^{n-i}(\mathbb{R})$.
(4) Note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{1},\ldots,x_{n+1})\neq(0,\ldots,0)$. Therefore, we can define a continuous function $Z\to\mathbb{P}^{n}(\mathbb{R})$ as follows:
$[x_{0}:\cdots:x_{n+1}]\mapsto[x_{1}:\cdots:x_{n+1}].$
As $d$ is odd, this is a bijection, and thus, a homeomorphism. So $Z\approx\mathbb{P}^{n}(\mathbb{R})$.
(5) First, note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{n},x_{n+1})\neq(0,0)$. Therefore, we can define a continuous map $Z\to\mathbb{P}^{i}(\mathbb{R})$ as follows:
$[x_{0}:\cdots:x_{n+1}]\mapsto[x_{n}:x_{n+1}].$
Then, this image has $\frac{d}{2}$ connected components
$E_{m}:=\left\{[\cos\theta:\sin\theta]\bigg{|}\frac{(4m-1)\pi}{2d}\leq\theta\leq\frac{(4m+1)\pi}{2d}\right\}$
for $10\leq m\leq\frac{d}{2}-$ 1.
In fact, $[x_{n}:x_{n+1}]\in f(Z)$ if and only if $(x_{n}+\sqrt{-1}x_{n+1})^{d}+(x_{n}-\sqrt{-1}x_{n+1})^{d}\leq 0$, and
$(\cos\theta+\sqrt{-1}\sin\theta)^{d}+(\cos\theta-\sqrt{-1}\sin\theta)^{d}=2\cos d\theta.$
Also, for each $m$, we can define a continuous map $f^{-1}(E_{m})\to S^{n}$ defined as follows:
$[x_{0}:\cdots:x_{n-1}:\cos\theta:\sin\theta]\mapsto\frac{1}{\sqrt{x_{0}^{2}+\cdots x_{n-1}^{2}+\sin d\theta}}(x_{0},\ldots,x_{n-1},\sin d\theta)$
Note that this is well-defined as $d$ is even. As this is a bijection, it is also a homeomorphism, and thus $Z\approx\underbrace{S^{n}\sqcup\cdots\sqcup S^{n}}_{\frac{d}{2}}$. ∎
In particular, we determined the real locus of real forms for the curve cases:
Theorem 5.4.
  1. (1)
    Suppose $d$ is odd. Then, the real loci of
    $(X_{0}^{d}+X_{1}^{d}+X_{2}^{d}=0),\ (X_{0}^{d}+(X_{1}+\sqrt{-1}X_{2})^{d}+(X_{1}-\sqrt{-1}X_{2})^{d}=0)$
    are both homeomorphic to $\mathbb{P}^{n}(\mathbb{R})$.
  2. (2)
    Suppose $d$ is even. Then,
    • the real locus of $(X_{0}^{d}+X_{1}^{d}+X_{2}^{d}=0)$ is $\emptyset$.
    • the real locus of $(-X_{0}^{d}+X_{1}^{d}+X_{2}^{d}=0)$ is homeomorphic to $S^{1}$.
    • the real locus of $(X_{0}^{d}+(X_{1}+\sqrt{-1}X_{2})^{d}+(X_{1}-\sqrt{-1}X_{2})^{d}=0)$ is homeomorphic to $\underbrace{S^{1}\sqcup\cdots\sqcup S^{1}}_{\frac{d}{2}}$.
Remark 5.5.
By considering real loci, we see that $F^{2}_{4}$ has at least $4$ non-isomorphic real forms
  • $(X_{0}^{4}+X_{1}^{4}+X_{2}^{4}+X_{3}^{4}=0)$,
  • $(-X_{0}^{4}+X_{1}^{4}+X_{2}^{4}+X_{3}^{4}=0)$,
  • $(-X_{0}^{4}-X_{1}^{4}+X_{2}^{4}+X_{3}^{4}=0)$, and
  • $(X_{0}^{4}+X_{1}^{4}+(X_{2}+\sqrt{-1}X_{3})^{4}+(X_{2}-\sqrt{-1}X_{3})^{4}=0)$.
Note that $F^{2}_{4}$ is a K3 surface and thus has only finitely many real forms up to isomorphisms by [CF20]. However, unlike the other cases $d\geq 3$, the automorphism group of $F^{2}_{4}$ is still unknown, and it is known to be a very difficult problem to determine the automorphism group of $F^{2}_{4}$. For this reason, it is also a very challenging problem to determine all real forms of $F^{2}_{4}$ or to see at least if there is a real form of $F^{2}_{4}$ other than these four or not.
Remark 5.6.
Let $Q^{r,s}$ be a hypersurface defined as follows:
$Q^{r,s}:=(X_{0}^{2}+\cdots+X_{r-1}^{2}-X_{r}^{2}-\cdots-X_{r+s-1}^{2}=0)\subset\mathbb{P}^{r+s-1}.$
Then, $Q^{r,s}$ with $r+s=n+2$ is a real form of $F^{n}_{2}$. Also, by considering real loci, we see that $Q^{r,s}$ and $Q^{t,u}$ with $r+s=t+u=n+2$ are $\mathbb{R}$-isomorphic if and only if $r=t$ or $r=u$.
However, there exists a real form of $F^{n}_{2}$ other than these forms. For example, by [Rus02], any real forms of $F^{2}_{2}$ are exactly $\mathbb{R}$-isomorphic to $Q^{4,0}$, $Q^{3,1}$, $Q^{2.2}$, or $Q^{3,0}\times Q^{2,1}$.

### v2

@@
 Abstract.
-In this paper, we compute the number of real forms of Fermat hypersurfaces for degree $d\geq 3$ except the degree $4$ surface case, and give explicit descriptions of them.
+In this paper, we compute the number of real forms of Fermat hypersurfaces for degree $d$ except the degree $4$ surface case, and give explicit descriptions of them.
 1. Introduction
@@
 Suppose $(n,d)\neq(2,4)$ and $d\geq 3$. Then, the number of real forms of $F^{n}_{d,\mathbb{C}}$ is
-$\begin{cases}\left\lfloor\frac{n}{2}\right\rfloor+2&(d\equiv 1\pmod{2})\\ \frac{(n+3)(n+5)}{8}&(d\equiv 0,\ n\equiv 1\pmod{2})\\ \frac{(n+4)(n+6)}{8}&(d\equiv 0,\ n\equiv 0\pmod{2})\end{cases}$
+$\begin{dcases}\left\lfloor\frac{n}{2}\right\rfloor+2&(d\equiv 1\pmod{2})\\ \frac{(n+3)(n+5)}{8}&(d\equiv 0,\ n\equiv 1\pmod{2})\\ \frac{(n+4)(n+6)}{8}+1&(d\equiv 0,\ n\equiv 0\pmod{2})\end{dcases}$.
+Also, the number of real forms of $F^{n}_{2,\mathbb{C}}$ is
+$\begin{dcases}\frac{n+1}{2}+1&(n\equiv 1\pmod{2})\\ \frac{n}{2}+3&(n\equiv 0\pmod{2})\end{dcases}$.
 We prove this theorem by computing $H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})$.
-In Section 3, we deal with the case $(n,d)\neq(1,3),(2,4)$, $d\geq 3$, and in Section 4, we deal with the case $(n,d)=(1,3)$.
-In both cases, we have already known the automorphism group $\operatorname{\operatorname{Aut}}F^{n}_{d}$, as the former case is due to Shioda [Shi87], and the latter is due to the fact that $F^{1}_{3}$ is elliptic.
+In Section 3, we deal with the case $(n,d)\neq(1,3),(2,4)$, $d\geq 3$, in Section 4, we deal with the case $(n,d)=(1,3)$, and in Section 5, we deal with the case $d=2$.
+In every case, we have already known the automorphism group $\operatorname{\operatorname{Aut}}F^{n}_{d}$, as the first case is due to Shioda [Shi87], the second is due to the fact that $F^{1}_{3}$ is elliptic, and in the third case, $\operatorname{\operatorname{Aut}}F^{n}_{2}\cong\operatorname{PO}(n+2)$.
 So we will check the condition for an automorphism of $F^{n}_{d}$ to be a cocycle, compute a conjugation of a cocycle by an automorphism, and get this result.
-We also give explicit descriptions of these real forms for the curve case in Theroem 5.1, and for the surface case in Theorem 5.2 in the way we can generalize it to higher dimensions, and give a topological description of a real locus of some real forms (in Theorem 5.3, and in particular, Theorem 5.4 for the curve case).
+We also give explicit descriptions of these real forms for the curve case in Theorem 6.2, and for the surface case in Theorem 6.3 in the way we can generalize it to higher dimensions, and give a topological description of a real locus of some real forms (in Theorem 6.4, and in particular, Theorem 6.5 for the curve case).
 We also show that $F_{4}^{2}$ has at least $4$ non-isomorphic real forms in Remark 6.6.
+This paper is a refined version of my master thesis [S21] (in Japanese, unpublished) under Professor Keiji Oguiso, together with a correction of some mistakes in the preliminary version, which was kindly pointed out by Professor János Kollár.
 2. Preliminary about real forms
@@
   4. (4)
-    Real structures $\rho$, $\rho^{\prime}$ on $X$ are equivalent if there exists a $\varphi\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ such that $\rho^{\prime}=\varphi^{-1}\circ\rho\circ\varphi$.
+    Real structures $\rho$, $\rho^{\prime}$ on $X$ are equivalent if there exists a $\phi\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ such that $\rho^{\prime}=\phi^{-1}\circ\rho\circ\phi$.
 Theorem 2.2 ([BS64]).
-Let $X$ be a complex quasi-projective variety with a real structure $\rho$ on $X$. Then, there are bijections between
+Let $X$ be a complex quasi-projective variety with a real structure $\rho$. Then, there are bijections between
   • the set of $\mathbb{R}$-isomorphism classes of real forms of $X$,
@@
   • $\alpha\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ is a cocycle if and only if $(\alpha\circ\rho)^{2}=\operatorname{id}_{X}$.
-  • cocycles $\alpha,\beta\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ are equivalent if and only if there exists $\varphi\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ such that $\beta\circ\rho=\varphi^{-1}\circ(\alpha\circ\rho)\circ\varphi$, and we denote it by $\alpha\sim\beta$.
+  • cocycles $\alpha,\beta\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ are equivalent if and only if there exists $\phi\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ such that $\beta\circ\rho=\phi^{-1}\circ(\alpha\circ\rho)\circ\phi$, and we denote it by $\alpha\sim\beta$.
 3. The number of real forms of Fermat hypersurfaces
-Before thinking about real forms of Fermat hypersurfaces, we want to think about automorphisms of them.
-Clearly, there are two types of automorphisms when the base field is algebraically closed and of characteristic 00.
-An automorphism of the first type is the one with which we can identify $\sigma\in S_{n+2}$ defined as follows:
+In order to study real forms of Fermat hypersurfaces, we first recall automorphisms of them.
+Clearly, there are two types of automorphisms when the base field is algebraically closed and of characteristic 0.
+An automorphism of the first type is the one with which we can identify $\sigma\in\mathfrak{S}_{n+2}$ defined as follows:
 $F^{n}_{d}\to F^{n}_{d};[X_{0}:\cdots:X_{n+1}]\mapsto[X_{\sigma(0)}:\cdots:X_{\sigma(n+1)}]$
 with $\sigma$ viewed as a permutation of $\{0,\ldots,n+1\}$.
-An automorphism of the second type is the one denoted by $g(i,j)$ for $i,j\in\mathbb{Z}$, $10\leq i\leq n+$ 1 and defined as follows:
+An automorphism of the second type is the one denoted by $g(i,j)$ for $i,j\in\mathbb{Z}$, $0\leq i\leq n+1$ and defined as follows:
 $F^{n}_{d}\to F^{n}_{d};[X_{0}:\cdots X_{n+1}]\mapsto[X_{0}:\cdots:\xi^{j}X_{i}:\cdots:X_{n+1}]$
-with $\xi=\xi_{d}$ a primitive $d$-th root of unity.
-Let $G^{n}_{d}$ be the subgroup of $\operatorname{\operatorname{Aut}}F^{n}_{d}$ generated by $g(i,j)$ for all $i,j\in\mathbb{Z}$. Then, the following result is known.
+with $\xi=\xi_{d}$ a primitive $d$-th root of unity. Also, we denote the product $g(0,j_{0})\cdot\cdots\cdot g(n,j_{n})$ by $D\left(j_{0},\ldots,j_{n}\right)$.
+Let $G^{n}_{d}$ be the subgroup of $\operatorname{\operatorname{Aut}}F^{n}_{d}$ generated by $g(i,j)$ for all $i,j\in\mathbb{Z}$. Then, the following result is known:
 Theorem 3.1 ([Shi87]).
 Let $k$ be an algebraically closed field with $\operatorname{char}k=0$, and suppose $(n,d)\neq(1,3),(2,4)$ and $d\geq 3$. Then,
-$\operatorname{\operatorname{Aut}}F^{n}_{d,k}=G^{n}_{d}\cdot S_{n+2}$
-where normal-$\cdot$ denotes a semidirect product.
-In particular, $\operatorname{\operatorname{Aut}}F^{n}_{d,\mathbb{C}}=G^{n}_{d}\cdot S_{n+2}$.
-Though this theorem is proved only in the curve case in that paper, the same argument works in higher dimensions as we can show that every automorphism is linear.
+$\operatorname{\operatorname{Aut}}F^{n}_{d,k}=G^{n}_{d}\cdot\mathfrak{S}_{n+2}$
+where $\cdot$ denotes a semidirect product.
+In particular, $\operatorname{\operatorname{Aut}}F^{n}_{d,\mathbb{C}}=G^{n}_{d}\cdot\mathfrak{S}_{n+2}$.
+Though this theorem is proved only in the curve case in that paper, the same argument works in higher dimensions, as we can show that every automorphism is linear.
 From now on, we assume $(n,d)\neq(1,3),(2,4)$, $d\geq 3$, and the base field is $\mathbb{C}$ unless otherwise stated.
-Let $\tilde{\rho}\colon\mathbb{P}^{n+1}\to\mathbb{P}^{n+1}$ be the complex conjugation, i.e., the $\mathbb{R}$-morphism defined by $\mathbb{C}[X]\to\mathbb{C}[X]$ that sends $X_{i}$ to $X_{i}$ for all $i$ and $c\in\mathbb{C}$ to $\bar{c}$.
+Let $\tilde{\rho}\colon\mathbb{P}^{n+1}\to\mathbb{P}^{n+1}$ be the complex conjugation, i.e., the $\mathbb{R}$-morphism defined by $\mathbb{C}[X]\to\mathbb{C}[X]$ that sends $X_{i}$ to $X_{i}$ for all $i$ and $c\in\mathbb{C}$ to $\overline{c}$.
 This is a real structure of $\mathbb{P}^{n+1}$, and it induces the real structure $\rho\colon F^{n}_{d}\to F^{n}_{d}$ on $F^{n}_{d}$.
 Remark 3.2.
-By straightforward calculations, for $i,j\in\mathbb{Z}$, $10\leq i\leq n+$ 1, $\sigma\in S_{n+2}$,
-  • $g(i,j)\circ\sigma=\sigma\circ g(\sigma^{-1}(i),j)$.
-  • $\sigma\circ\rho=\rho\circ\sigma$.
-  • $g(i,j)\circ\rho=\rho\circ g(i,d-j)=\rho\circ g(i,j)^{-1}$.
-The following result is the main step to prove the main theorem.
+The following relations hold for $D\left(j_{0},\ldots,j_{n+1}\right)$, $\sigma\in\mathfrak{S}_{n+2}$, and $\rho$.
+$\sigma\cdot D\left(j_{0},\ldots,j_{n+1}\right)\cdot\sigma^{-1}=D\left(j_{\sigma(0)},\ldots,j_{\sigma(n+1)}\right)$
+$\sigma\cdot\rho=\rho\cdot\sigma$
+$D\left(j_{0},\ldots,j_{n+1}\right)\cdot\rho=\rho\cdot D\left(-j_{0},\ldots,-j_{n+1}\right)=\rho\cdot D\left(j_{0},\ldots,j_{n+1}\right)^{-1}.$
+The following result is the main step in proving the main theorem.
 Proposition 3.3.
+Let $(n,d)\neq(1,3),(2,4)$, and $d\geq 3$.
   1. (1)
-    $\alpha=\mu\sigma\in\operatorname{\operatorname{Aut}}F^{n}_{d}$, with $\mu=g(0,p_{0})\cdots g(n+1,p_{n+1})\in G_{d}^{n}$, $\sigma\in S_{n+2}$, is a cocycle if and only if the following two conditions hold:
+    $\alpha=\mu\sigma\in\operatorname{\operatorname{Aut}}F^{n}_{d}$, with $\mu=D\left(p_{0},\ldots,p_{n+1}\right)\in G_{d}^{n}$, $\sigma\in\mathfrak{S}_{n+2}$, is a cocycle if and only if the following conditions are satisfied:
     • $\sigma$ is of type $1^{l}\cdot 2^{m}$ with $l+2m=n+2$.
-    • $p_{i}\equiv p_{\sigma(i)}\pmod{d}$ if $i\in\operatorname{supp}\sigma$.
-    Here, $\operatorname{supp}\sigma$ is the support of a permutation $\sigma$, i.e., $\operatorname{supp}\sigma:=\left\{i|\sigma(i)\neq i\right\}$.
+    • $p_{i}\equiv p_{\sigma(i)}\pmod{d}$ for any $0\leq i\leq n+1$ if $\sigma$ has a fixed point, i.e., $\sigma(k)=k$ for some $k$.
+    • $p_{i}-p_{\sigma(i)}\equiv p_{j}-p_{\sigma(j)}\equiv 0,\dfrac{d}{2}\pmod{d}$ for any $0\leq i,j\leq n+1$ if $\sigma$ has no fixed point, i.e., $\sigma$ is of type $2^{\frac{n+2}{2}}$.
   2. (2)
-    Cocycles $\beta=\lambda\tau$, $\gamma$ with $\lambda\in G^{n}_{d}$ and $\tau\in S_{n+2}$ are equivalent if and only if there exist $\psi\in S_{n+2}$ and $q_{j}\in\mathbb{Z}$ for $10\leq j\leq n+$ 1 with $q_{j}\equiv q_{\sigma(j)}\pmod{2}$ for all $j$ and $q_{l}\equiv 0\pmod{2}$ for all $l\notin\operatorname{supp}\tau$ such that
-    $\gamma=\psi^{-1}\cdot\left(g(0,q_{0})\cdots g(n+1,q_{n+1})\beta\right)\cdot\psi.$
+    Cocycles $\beta=\lambda\tau$ with $\lambda\in G^{n}_{d}$ and $\tau\in\mathfrak{S}_{n+2}$ and $\gamma$ are equivalent if and only if there exist $\psi\in\mathfrak{S}_{n+2}$ and $r_{j}\in\mathbb{Z}$ for $0\leq j\leq n+1$ with $r_{j}=r_{\sigma(j)}$ for all $j$ and $r_{l}\equiv 0\pmod{2}$ for all $l\notin\operatorname{supp}\tau$ such that
+    $\gamma=\psi^{-1}\cdot\left(D\left(r_{0},\ldots,r_{n+1}\right)\cdot\beta\right)\cdot\psi.$
+    Here, $\operatorname{supp}\tau$ is the support of a permutation $\tau$, i.e., $\operatorname{supp}\tau:=\left\{i|\tau(i)\neq i\right\}$.
 Proof.
 (1) By definition, $\alpha$ is a cocycle if and only if $(\mu\sigma\rho)^{2}=\operatorname{id}$. Also,
-$(\mu\sigma\rho)^{2}$ $=\mu\sigma\mu^{-1}\sigma$
-$=g(0,{p^{\prime}}_{0})\cdots g(n+1,{p^{\prime}}_{n+1})\cdot\sigma^{2}$
-where
-${p^{\prime}}_{i}=\begin{cases}0&(i\notin\operatorname{supp}\sigma)\\ p_{i}-p_{\sigma(i)}&(i\in\operatorname{supp}\sigma)\end{cases}$
-Therefore, $(\mu\sigma\rho)^{2}=\operatorname{id}$ is equivalent to $\sigma^{2}=\operatorname{id}$ and ${p^{\prime}}_{i}\equiv 0\pmod{d}$, and the result follows.
-(2) Suppose $\beta=\lambda\tau$ be a cocycle and $\varphi=\nu\psi\in\operatorname{\operatorname{Aut}}F^{n}_{d}$ with $\lambda,\nu=g(0,r_{0})\cdots g(n+1,r_{n+1})\in G^{n}_{d},\ \tau,\psi\in S_{n+2}$.
+$(\mu\sigma\rho)^{2}$ $=\mu\sigma\rho\mu\sigma\rho=\mu\sigma\mu^{-1}\sigma$
+$=D\left(p_{0},\ldots,p_{n+1}\right)\cdot D\left(-p_{\sigma(0)},\ldots,-p_{\sigma(n+1)}\right)\cdot\sigma^{2}$
+$=D\left(p_{0}-p_{\sigma(0)},\ldots,p_{n+1}-p_{\sigma(n+1)}\right)\cdot\sigma^{2}$
+Therefore, $(\mu\sigma\rho)^{2}=\operatorname{id}$ is equivalent to $(\mu\sigma\rho)^{2}=D\left(N,\ldots,N\right)$ for some $N\in\mathbb{Z}$.
+When $\sigma(i)=i$, as $p_{i}-p_{\sigma(i)}=0$, $(\mu\sigma\rho)^{2}=\operatorname{id}$ is equivalent to $\sigma^{2}=\operatorname{id}$ and $p_{j}-p_{\sigma(j)}\equiv 0\pmod{d}$ for any $j$.
+When $\sigma$ has no fixed point, $(\mu\sigma\rho)^{2}=\operatorname{id}$ is equivalent to $\sigma^{2}=\operatorname{id}$ and $p_{0}-p_{\sigma(0)}\equiv\cdots\equiv p_{n+1}-p_{\sigma(n+1)}\pmod{d}$.
+In this case, for $0\leq i\leq n+1$,
+$p_{i}-p_{\sigma(i)}\equiv p_{\sigma(i)}-p_{\sigma^{2}(i)}\equiv p_{\sigma(i)}-p_{i}\pmod{d}$
+as $\sigma^{2}=\operatorname{id}$. Therefore, in this case, $p_{i}-p_{\sigma(i)}\equiv 0,\frac{d}{2}$ as $\sigma(i)\neq i$, and this is exactly what we want to show.
+(2) Suppose $\beta=\lambda\tau$ be a cocycle and $\phi=\nu\psi\in\operatorname{\operatorname{Aut}}F^{n}_{d}$ with $\lambda,\nu=D\left(q_{0},\ldots,q_{n+1}\right)\in G^{n}_{d},\ \tau,\psi\in\mathfrak{S}_{n+2}$.
 Then,
-$\varphi^{-1}\circ\beta\rho\circ\varphi$
-$=$ $(\psi^{-1}\nu^{-1})\cdot(\lambda\tau\rho)\cdot(\nu\psi)$
-$=$ $\psi^{-1}\cdot(\nu^{-1}\lambda\tau\nu^{-1})\cdot\psi\rho$
-$=$ $\psi^{-1}\cdot(\nu^{-1}\lambda\cdot g(\tau(0),d-r_{0})\cdots g(\tau(n+1),d-r_{n+1}))\cdot\tau\psi\rho$
-$=$ $\psi\cdot(g(0,s_{0})\cdots g(n+1,s_{n+1})\cdot\lambda\tau)\cdot\psi\rho$
-where
-$s_{i}=\begin{cases}-2r_{i}&(i\notin\operatorname{supp}\tau)\\ -r_{i}-r_{\tau(i)}&(i\in\operatorname{supp}\tau)\end{cases}$
-As $r_{i}$ varies for all integers, the result follows. ∎
-Note that, as cocycles,
+$\phi^{-1}\cdot\beta\rho\cdot\phi$
+$=$ $\left(\psi^{-1}\nu^{-1}\right)\cdot\left(\lambda\tau\rho\right)\cdot\left(\nu\psi\right)$
+$=$ $\psi^{-1}\cdot\left(\nu^{-1}\lambda\tau\nu^{-1}\right)\cdot\psi\rho$
+$=$ $\psi^{-1}\cdot\left(\nu^{-1}\lambda\cdot D\left(-q_{\tau(0)},\ldots,-q_{\tau(n+1)}\right)\tau\right)\cdot\psi\rho$
+$=$ $\psi^{-1}\cdot\left(D\left(q_{0}+q_{\tau(0)},\ldots,q_{n+1}+q_{\tau(n+1)}\right)^{-1}\lambda\tau\right)\cdot\psi\rho$
+As $q_{i}$’s vary for all integers, the result follows. ∎
+Remark 3.4.
+By Proposition 3.3, any cocycles in $\operatorname{\operatorname{Aut}}F^{n}_{d}$ are equivalent to one of the following forms:
+$\begin{pmatrix}\xi^{a_{0}}&&\\ &\ddots&\\ &&\xi^{a_{n+1}}\end{pmatrix}\cdot\begin{pmatrix}I_{r}&&&&&\\ &0&1&&&\\ &1&0&&&\\ &&&\ddots&&\\ &&&&0&1\\ &&&&1&0\end{pmatrix}$
+with $I_{r}$ the identity matrix of size $r$, $a_{i}\in\mathbb{Z}$, and $0\leq r\leq n+2$. Throughtout this paper, blank entries in matrices mean 0. Also, as cocycles,
 $I_{r}=\xi^{m}I_{r}$
-for $m\in\mathbb{Z}$ by definition, and
-$\begin{pmatrix}1&\\ &I_{r}\end{pmatrix}\sim\begin{pmatrix}\xi^{2i}&\\ &I_{r}\end{pmatrix},\ \begin{pmatrix}I_{r}&&\\ &0&1\\ &1&0\end{pmatrix}\sim\begin{pmatrix}I_{r}&&\\ &0&\xi^{j}\\ &\xi^{j}&0\end{pmatrix}$
-for $i,j\in\mathbb{Z}$ by the proposition.
-By this proposition, we can describe $H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})$.
+for $m\in\mathbb{Z}$ by definition, and in $\operatorname{\operatorname{Aut}}\left(F_{d}^{n}\right)$
+$\begin{pmatrix}1&\\ &A\end{pmatrix}\sim\begin{pmatrix}\xi^{2i}&\\ &A\end{pmatrix},\ \begin{pmatrix}B&&\\ &0&1\\ &1&0\end{pmatrix}\sim\begin{pmatrix}B&&\\ &0&\xi^{j}\\ &\xi^{j}&0\end{pmatrix}$
+for $i,j\in\mathbb{Z}$ and any matrices $A,B$ by Proposition 3.3.
+Now, we can describe $H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})$.
 Theorem 3.5.
-Suppose $d$ is odd. Then, any cocycles are equivalent to exactly one of the following forms:
-[…]
+Suppose $d$ is odd. Then, any cocycles in $\operatorname{\operatorname{Aut}}F^{n}_{d}$ are equivalent to exactly one of the following forms:
+$H^{r}_{n}:=\begin{pmatrix}I_{r}&&&&&\\ &0&1&&&\\ &1&0&&&\\ &&&\ddots&&\\ &&&&0&1\\ &&&&1&0\end{pmatrix}\ (r\in\mathbb{Z},\ 0\leq r\leq n+2).$
 In particular, $F^{n}_{d}$ has exactly $\left\lfloor\frac{n}{2}\right\rfloor+2$ real forms.
@@
 As $d$ is odd,
-$\begin{pmatrix}1&\\ &I_{r}\end{pmatrix}\sim\begin{pmatrix}\xi^{i}&\\ &I_{r}\end{pmatrix}$
-for all $i$, and the former statement follows from Proposition 3.3. Also, as any two of $H^{r}_{n}$’s are not equivalent,
+$\begin{pmatrix}1&\\ &A\end{pmatrix}\sim\begin{pmatrix}\xi^{i}&\\ &A\end{pmatrix}$
+for all $i$ and matrices AA.
+Therefore, when $d$ is odd, cocycles $\mu\sigma$, $\lambda\tau$ with $\mu,\lambda\in G_{d}^{n}$ and $\sigma,\tau\in\mathfrak{S}_{n+2}$ are equivalent if and only if the types of $\sigma$ and $\tau$ agree by Proposition 3.3 and Remark 3.4.
+Therefore,
 $H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})=\left\{H^{r}_{n}|0\leq r\leq n+2\right\}$
-where we identify cocycles with their equivalent classes, and the latter statement follows. ∎
+where we identify cocycles with their equivalent classes, and the statement follows. ∎
 Theorem 3.6.
-Suppose $d$ is even. Then, any cocycles are equivalent to exactly one of the forms:
+Suppose $d$ is even. Then, any cocycles in $\operatorname{\operatorname{Aut}}F^{n}_{d}$ are equivalent to exactly one of the forms:
 $K^{s,t}_{n}:=\begin{pmatrix}\xi I_{s}&&&&&&\\ &I_{t}&&&&&\\ &&0&1&&&\\ &&1&0&&&\\ &&&&\ddots&&\\ &&&&&0&1\\ &&&&&1&0\end{pmatrix}\ (s,t\in\mathbb{Z}_{\geq 0},\ s+t\leq n+2)$
-Also, $F^{n}_{d}$ has exactly $\frac{(n+3)(n+5)}{8}$ real forms for $n$ odd, and $\frac{(n+4)(n+6)}{8}$ real forms for $n$ even.
-In particular, if $d\geq 3$ is even, then $F^{1}_{d}$ has exactly $3$ real forms and $F^{2}_{d}$ ($d\not=4$) has $6$ real forms.
+$L_{n}:=\begin{pmatrix}0&1&&&\\ -1&0&&&\\ &&\ddots&&\\ &&&0&1\\ &&&-1&0\end{pmatrix}\ (n\equiv 0\pmod{2})$
+Also, $F^{n}_{d}$ has exactly $\frac{(n+3)(n+5)}{8}$ real forms for $n$ odd, and $\frac{(n+4)(n+6)}{8}+1$ real forms for $n$ even.
+In particular, if $d\geq 3$ is even, then $F^{1}_{d}$ has exactly $3$ real forms and $F^{2}_{d}$ ($d\not=4$) has $7$ real forms.
 Proof.
-The former statement is clear from proposition 3.3.
-By the proposition, for $s,s^{\prime},t,t^{\prime}\in\mathbb{Z}$ with $s+t=s^{\prime}+t^{\prime}$, $K^{s,t}_{n}\sim K^{s^{\prime},t^{\prime}}$ if and only if there exists a permutation matrix $\sigma$ and $a_{1}$, $\ldots$, $a_{n+2}$ with $a_{1}\equiv\cdots\equiv a_{m}\pmod{2}$ such that
-$K^{s,t}_{n}\cdot\left(\sigma^{-1}K^{s^{\prime},t^{\prime}}_{n}\sigma\right)^{-1}=\begin{pmatrix}\xi^{a_{1}}&&\\ &\ddots&\\ &&\xi^{a_{n+2}}\end{pmatrix}.$
-This condition is equivalent to $s=s^{\prime}$ or $s=t^{\prime}$.
+By Proposition 3.3 and Remark 3.4, any cocycles are equivalent to $L_{n}$ or
+$d_{r}(a_{1},\ldots,a_{r}):=\begin{pmatrix}\xi^{a_{1}}&&&&&&&\\ &\ddots&&&&&&\\ &&\xi^{a_{r}}&&&&&\\ &&&0&1&&&\\ &&&1&0&&&\\ &&&&&\ddots&&\\ &&&&&&0&1\\ &&&&&&1&0\end{pmatrix}$
+with $a_{i}=0,1$ for each $i$.
+Then, by Proposition 3.3, $\alpha:=d_{r}(a_{1},\ldots,a_{r})$ is equivalent to $\beta:=d_{r}(b_{1},\ldots,b_{r})$ if and only if $\beta\sigma^{-1}\alpha^{-1}\sigma=d_{r}(c_{1},\ldots,c_{r})$ with $\sigma\in\mathfrak{S}_{n+2}$ and $c_{i}\equiv c_{j}\pmod{2}$ for all $i,j$, which is equivalent to say that $n(\alpha)=n(\beta)$ or $n(\alpha)+n(\beta)=r$ where $n(\alpha)=\#\left\{i|a_{i}=1\right\}$.
+Also, $d_{r}(a_{1},\ldots,a_{r})$ and $d_{s}(b_{1},\ldots,b_{s})$ are not equivalent if $r\neq s$ by Propositon 3.3.
 Therefore,
-[…]
-where we identify cocycles with their equivalent classes. From this, when $n=2k-1$ for $k\in\mathbb{Z}$, the number of real forms is
+$H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})\setminus\{L_{n}\}=\left\{K^{s,t}_{n}|s+t\leq n+2,\ 0\leq s\leq t\right\}$
+where we identify cocycles with their equivalent classes. From this, when n=2k1n=2k-1 for $k\in\mathbb{Z}$, the number of real forms is
 $\sum_{i=0}^{k}(i+1)=\frac{(k+1)(k+2)}{2}=\frac{(n+3)(n+5)}{8},$
 and when $n=2k$ for $k\in\mathbb{Z}$, the number of real forms is
-$\sum_{i=0}^{k+1}(i+1)=\frac{(k+2)(k+3)}{2}=\frac{(n+4)(n+6)}{8},$
+$1+\sum_{i=0}^{k+1}(i+1)=1+\frac{(k+2)(k+3)}{2}=\frac{(n+4)(n+6)}{8}+1,$
 so the result follows. ∎
-We omit subindexes of $H^{r}_{n}$ or $K^{s,t}_{n}$ if it is clear from the context.
 4. The elliptic curve case
@@
 $X=\frac{-12X_{2}}{X_{0}+X_{1}},\ Y=\frac{36(X_{0}-X_{1})}{X_{0}+X_{1}}$
-gives an elliptic curve $Y^{2}=X^{3}-432$ with $j$-invariant 00, and this can be identified with $E:=\mathbb{C}/(\mathbb{Z}+\zeta\mathbb{Z})$ where $\zeta$ is a primitive $6$-th root of unity, which we fix in this section.
+gives an elliptic curve $Y^{2}=X^{3}-432$ with $j$-invariant 0, and this can be identified with $E:=\mathbb{C}/(\mathbb{Z}+\zeta\mathbb{Z})$ where $\zeta$ is a primitive $6$-th root of unity, which we fix in this section.
 Then, any automorphisms of $E$ can be expressed by
@@
 [Sil09]).
-Note that the real structure $\rho\colon F^{1}_{3}\to F^{1}_{3}$ corresponds to $E\to E$ that maps $z$ to $\bar{z}$ where we identify $z\in\mathbb{C}$ with the image of $z$ by the canonical surjection $\mathbb{C}\to\mathbb{C}/(\mathbb{Z}+\zeta\mathbb{Z})$.
+Note that the real structure $\rho\colon F^{1}_{3}\to F^{1}_{3}$ corresponds to $E\to E$ that maps $z$ to $\overline{z}$ where we identify $z\in\mathbb{C}$ with the image of $z$ by the canonical surjection $\mathbb{C}\to\mathbb{C}/(\mathbb{Z}+\zeta\mathbb{Z})$.
 We also denote it by $\rho\colon E\to E$.
@@
 (1) By definition, $\alpha$ is a cocycle if and only if $(\alpha\circ\rho)^{2}=\operatorname{id}$.
-$(\alpha\circ\rho)^{2}(z)$ $=\zeta^{i}(\overline{\zeta^{i}\bar{z}+a})+a$
-$=z+(\zeta^{i}\bar{a}+a)$
+$(\alpha\circ\rho)^{2}(z)$ $=\zeta^{i}(\overline{\zeta^{i}\overline{z}+a})+a$
+$=z+(\zeta^{i}\overline{a}+a)$
 Therefore, the result follows.
@@
 $=$ $\tau^{-1}(\zeta^{i}(\overline{\zeta^{j}z+b})+a)$
-$=$ $\zeta^{i-2j}\bar{z}+\zeta^{-j}a-\zeta^{j}b+\zeta^{i-j}\bar{b}$
+$=$ $\zeta^{i-2j}\overline{z}+\zeta^{-j}a-\zeta^{j}b+\zeta^{i-j}\overline{b}$
 First, $z\mapsto z$ and $z\mapsto-z$ are not equivalent to each other by this calculation.
@@
 $j=0,\ b=\frac{1+\sqrt{-1}}{2}\xi^{i}x,$
-we see that $z\to\zeta^{i}z+a$ is equivalent to $z\to\zeta^{i}z$, so the result follows. ∎
+we see that $z\to\zeta^{i}z+a$ is equivalent to $z\to\zeta^{i}z$, so, by setting $a=b=0$ in the above calculation, we get the result. ∎
 Note that the negation map in the above corresponds to a linear automorphism
 $F^{1}_{3}\to F^{1}_{3};\ [X_{0}:X_{1}:X_{2}]\mapsto[X_{1}:X_{0}:X_{2}]$
-by taking $[1:-1:0]$ as a unit. From this, we immediately get the following result:
+by taking [1:1:0][1:-1:0] as a unit. From this, we immediately get the following result:
 Theorem 4.2.
@@
 In particular, $F^{1}_{3}$ has exactly $2$ real forms.
+5. Degree $2$ cases
+(This section is based on the advice from Professor János Kollár. )
+In this section, we consider the number and explicit descriptions of the real forms of $X_{n}:=F^{n}_{2}=\left(X_{0}^{2}+\cdots+X_{n+1}^{2}=0\right)\subset\mathbb{P}^{n+1}$.
+Let $Q_{r,s}\subset\mathbb{P}^{r+s-1}$ be a quadratic hypersurface defined as follows:
+$Q_{r,s}:=\left(X_{0}^{2}+\cdots+X_{r-1}^{2}-X_{r}^{2}-\cdots X_{r+s-1}^{2}=0\right)\subset\mathbb{P}^{r+s-1}.$
+Then, $F^{1}_{2}$ is isomorphic to $\mathbb{P}^{1}$, and has exactly two real forms $Q_{2,0}$ and $Q_{1,1}$ up to isomorphism.
+Also, it is known that $F^{2}_{2}$ has exactly four real forms $Q_{4,0}$, $Q_{3,1}$, $Q_{2,2}$, and $Q_{3,0}\times Q_{2,1}$ up to isomorphisms (cf.
+[Rus02]).
+Therefore, we consider the case where $n\geq 3$.
+Lemma 5.1.
+When $n\geq 3$, any real forms of $X_{n}$ can be embedded into real forms of $\mathbb{P}^{n+1}$.
+Proof.
+Let $H$ be a hyperplane section of $X_{n}$. Then, $|H|$ induces an embedding $\varphi_{|H|}\colon X_{n}\hookrightarrow\mathbb{P}^{n+1}$.
+Considering the automorphism group of $X_{n}$, we see that $X_{n}$ is also a smooth hypersurface of degree $2$ in $\mathbb{P}^{n+1}$ in this embedding.
+Then, every real structure of $X_{n}$ preserves $|H|$ as every automorphism of $X_{n}$ is linear.
+Therefore, by considering the quotient by a real structure, we see that every real form of $X_{n}$ can be embedded into some real form of $\mathbb{P}^{n+1}$.
+∎
+Theorem 5.2.
+When $n$ is odd, any real forms of $F^{2}_{n}$ are isomorphic to $Q_{r,s}$ for some rrss with $r\geq s,r+s=n+2$.
+When $n$ is even, any real forms of $F^{2}_{n}$ are isomorphic to $Q_{r,s}$ for some rrss with $r\geq s,r+s=n+2$, or the real form corresponding to the cocycle $L_{n}$ in $\operatorname{\operatorname{Aut}}X_{n}$.
+In particular, when $n$ is even, $F^{2}_{n}$ has exactly $\frac{n}{2}+3$ real forms, and when $n$ is odd, $F^{2}_{n}$ has exactly $\frac{n+1}{2}+1$ real forms.
+Proof.
+By [Ben16], $P^{n+1}$ has exactly two real forms up to isomorphisms when $n$ is even, which correspond to the following two real structures:
+$\tilde{\rho}$ $\colon\mathbb{P}^{n+1}\to\mathbb{P}^{n+1};\ [z_{0}:\cdots:z_{n+1}]\mapsto\left[\overline{z_{0}}:\cdots:\overline{z_{n+1}}\right]$
+$\tilde{\rho^{\prime}}$ $\colon\mathbb{P}^{n+1}\to\mathbb{P}^{n+1};\ [z_{0}:\cdots:z_{n+1}]\mapsto\left[-\overline{z_{1}}:\overline{z_{0}}:\cdots:\overline{z_{n+1}}:-\overline{z_{n}}\right].$
+Here, we denote by $\widetilde{\mathbb{P}^{n+1}}$ the real form of $\mathbb{P}^{n+1}$ defined by the real structure $\tilde{\rho^{\prime}}$.
+Also, when $n$ is odd, $\mathbb{P}^{n+1}$ has exactly one real form up to isomorphisms, which corresponds to above $\tilde{\rho}$.
+Therefore, by Lemma 5.1, we enough to consider the case where a real form of $F^{2}_{n}$ is embedded into these two real forms of $\mathbb{P}^{n+1}$.
+First, suppose a real form $Y$ of $X_{n}$ is embedded into the real form of $\mathbb{P}^{n+1}$ corresponding to the real strucure $\tilde{\rho}$, i.e., $\mathbb{P}^{n+1}_{\mathbb{R}}$.
+Then, $Y$ is a hypersurface of degree $2$ of $\mathbb{P}^{n+1}_{\mathbb{R}}$ by the proof of Lemma 5.1.
+As $F^{n}_{2}$ is defined by a non-degenerated quadratic form, $Y$ is isomorphic to $Q_{r,s}$ for some $r,s$ with $r\geq s$ by considering a diagonalization of symmetric matrix.
+Second, suppose a real form $Y$ of $X_{n}$ is embedded into the real form of $\mathbb{P}^{n+1}$ corresponding to the real strucure $\tilde{\rho^{\prime}}$, i.e., $\widetilde{\mathbb{P}^{n+1}}$.
+Then, $\widetilde{\mathbb{P}^{n+1}}$ contains $\widetilde{\mathbb{P}^{1}}$, and as $F^{n}_{2}$ is defined by a non-degenerated quadratic form, we can take $L\subset\widetilde{\mathbb{P}^{n+1}}$ which is isomorphic to $\widetilde{\mathbb{P}^{1}}$ so that it is not contained in $Y$.
+We may assume that $L=\left(X_{2}=\cdots=X_{n+1}=0\right)\subset\widetilde{\mathbb{P}^{n+1}}$.
+Also, we may assume that $Y\cap L=\left(X_{0}^{2}+X_{1}^{2}=0\right)\subset L$ as hypersurfaces in $\widetilde{\mathbb{P}^{1}}$ defined by $\left(X_{0}^{2}+X_{1}^{2}=0\right)$ and $\left(X_{0}^{2}-X_{1}^{2}=0\right)$ is isomorphic.
+Then, by considering orthogonal complement, we can set the coordinate of $\widetilde{\mathbb{P}^{n+1}}$ so that $Y$ is defined by
+$X_{0}^{2}+X_{1}^{2}+q\left(X_{2},\ldots,X_{n+1}\right)=0$
+where $q$ is a non-degenerated quadratic form. Therefore, by recurring this procedure, we get the unique real form
+$\left(X_{0}^{2}+\cdots X_{n+1}^{2}=0\right)\subset\widetilde{\mathbb{P}^{n+1}}$
+of $X_{n}$ embedded in $\widetilde{\mathbb{P}^{n+1}}$, and it is clear that this real form correspond the cocycle $L_{n}$ in $\operatorname{\operatorname{Aut}}X_{n}$.
+By considering the real loci of $Q_{r,s}$, we see that $Q_{r,s}$’s with $r\geq s$ are not isomorphic to each other.
+Also, a real form $Q_{r,s}$ of $X_{n}$ corresponds to a cocycle $K^{s,r}_{2}$, and when $n$ is even, this cannot be equivalent to $L_{n}$.
+In fact, as matrices,
+$\begin{pmatrix}-I_{s}&\\ &I_{r}\end{pmatrix}^{2}=I_{r+s}$
+whereas
+$\begin{pmatrix}0&1&&&\\ -1&0&&&\\ &&\ddots&&\\ &&&0&1\\ &&&-1&0\end{pmatrix}^{2}=-I_{n+2}.$
+Therefore, we see that the number of real forms of $X_{n}$ is $\frac{n}{2}+3$ when $n$ is even, and $\frac{n+1}{2}+1$ when $n$ is odd. ∎
 6. Explicit description of real forms of Fermat hypersurfaces
-First, we consider the curve case $F^{1}_{d}=(X_{0}^{d}+X_{1}^{d}+X_{2}^{d}=0)$.
+In this section, we consider the explicit description of real forms corresponding to $H^{r}_{n}$, $K^{s,t}_{n}$, and $L_{n}$.
+We omit subindexes of $H^{r}_{n}$ or $K^{s,t}_{n}$ if it is clear from the context.
+First, we consider the real form corresponding to cocycles $L_{n}$.
+As we see in Theorem 3.5, Theorem 3.6, and Theorem 4.2, cocycles of this type appear only when $n$ and $d$ are even.
+Let $n=2k$.
+This is a real form induced by the real form $\widetilde{\mathbb{P}^{n+1}}$ of $\mathbb{P}^{n+1}$, i.e., corresponding to the real structure $\tilde{\rho^{\prime}}$.
+Therefore, the quotient $Z^{2k}_{d}$ of $F_{d,\mathbb{C}}^{2k}\subset\mathbb{P}_{\mathbb{C}}^{2k+1}$ with even $d$ by the restriction of $\rho^{\prime}$ is the real form corresponding to $L_{2k}$.
+Note that this real form can be embedded into $\widetilde{\mathbb{P}^{n+1}}$.
+Next, we consider the real form corresponding to cocycles $K$ or $H$. When considering these real forms, the following lemma is useful:
+Lemma 6.1.
+Let $\xi_{d}$ be a primitive $d$-th root of unity. Then,
+  1. (1)
+    The invariant ring of the $\mathbb{R}$-homomorphism $\alpha_{1}\colon\mathbb{C}[X]\to\mathbb{C}[X]$ that sends $X$ to $\xi_{d}X$ and $c\in\mathbb{C}$ to $\overline{c}$ is $\mathbb{R}\left[\xi_{2d}X\right]$, with $\xi_{2d}$ satisfying $\xi_{2d}^{2}=\xi_{d}$.
+  2. (2)
+    The invariant ring of the $\mathbb{R}$-homomorphism $\alpha_{2}\colon\mathbb{C}[X_{1},X_{2}]\to\mathbb{C}[X_{1},X_{2}]$ that sends $X_{1}$ to $X_{2}$, $X_{2}$ to $X_{1}$, and $c\in\mathbb{C}$ to $\overline{c}$ is $\mathbb{R}\left[X_{1}+X_{2},\left(1-\sqrt{-1}\right)\left(X_{1}-\sqrt{-1}X_{2}\right)\right]$.
+Proof.
+(1) For all $f\in\mathbb{C}[X]$, there is a unique polynomial $\mu(f)\in\mathbb{C}[X]$ such that
+$f(X)=\mu\left(\xi_{2d}X\right).$
+Then, as $\xi_{2d}X$ is invariant under $\alpha_{1}$, we see that $\alpha_{1}\left(f\right)=f$ if and only if $\mu(f)$ is a polynomial of real coefficients.
+This shows that the invariant ring with $\alpha_{1}$ is $\mathbb{R}\left[\xi_{2d}\right]$.
+(2) For all $g\in\mathbb{C}[X_{1},X_{2}]$, there is a unique polynomial $\nu(g)\in\mathbb{C}[X_{1},X_{2}]$ such that
+$g(X_{1},X_{2})=\nu(g)\left(X_{1}+X_{2},\left(1-\sqrt{-1}\right)\left(X_{1}-\sqrt{-1}X_{2}\right)\right).$
+Then, as $X_{1}+X_{2}$, $\left(1-\sqrt{-1}\right)\left(X_{1}-\sqrt{-1}X_{2}\right)$ is invariant under $\alpha_{2}$, we see that $\alpha_{2}(g)=g$ if and only if $\nu(g)$ is a polynomial with real coefficients.
+This shows that the invariant ring of $\alpha_{2}$ is $\mathbb{R}\left[X_{1}+X_{2},\left(1-\sqrt{-1}\right)\left(X_{1}-\sqrt{-1}X_{2}\right)\right]$.
+∎
+Then, we consider the curve case $F^{1}_{d}=(X_{0}^{d}+X_{1}^{d}+X_{2}^{d}=0)$.
 Theorem 6.2.
-Assume that $d\geq 3$. Then, the correspondence between cocycles and real forms of $F^{1}_{d}$ is given by the following table.
+Assume that $d\geq 3$. Then, the correspondence between cocycles and real forms of $F^{1}_{d}$ is given in the following table.
 cocycle real form
@@
 Next, we consider the real form corresponding to $K^{1,2}$.
-The invariant ring of the homomorphism $\alpha\colon\mathbb{C}[X_{0},X_{1},X_{2}]\to\mathbb{C}[X_{0},X_{1},X_{2}]$ that sends $c\in\mathbb{C}$ to $\bar{c}$, $X_{0}$ to $\xi_{d}X_{0}$, $X_{1}$ to $X_{1}$, and $X_{2}$ to $X_{2}$ is $\mathbb{R}[\xi_{2d}X_{0},X_{1},X_{2}]$ with $\xi_{2d}^{2}=\xi_{d}$.
-In fact, for all $f(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$, there exists a unique polynomial $\mu(f)(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$ such that
-$f(X_{0},X_{1},X_{2})=\mu(f)(\xi_{2d}X_{0},X_{1},X_{2}),$
-and $\alpha(f)=f$ if and only if $\mu(f)$ is a real coefficient polynomial, as $\xi_{2d}X_{0}$, $X_{1}$, $X_{2}$ are invariant under $\alpha$.
+By Lemma 6.1, the invariant ring of the homomorphism $\alpha\colon\mathbb{C}[X_{0},X_{1},X_{2}]\to\mathbb{C}[X_{0},X_{1},X_{2}]$ that sends $c\in\mathbb{C}$ to $\overline{c}$, $X_{0}$ to $\xi_{d}X_{0}$, $X_{1}$ to $X_{1}$, and $X_{2}$ to $X_{2}$ is $\mathbb{R}[\xi_{2d}X_{0},X_{1},X_{2}]$ with $\xi_{2d}^{2}=\xi_{d}$.
 Therefore, by setting
@@
 Finally, we consider the real form corresponding to $K^{1,2}$.
-The invariant ring of the homomorphism $\beta\colon\mathbb{C}[X_{0},X_{1},X_{2}]\to\mathbb{C}[X_{0},X_{1},X_{2}]$ that sends $c\in\mathbb{C}$ to $\bar{c}$, $X_{0}$ to $X_{0}$, $X_{1}$ to $X_{2}$ and $X_{2}$ to $X_{2}$ is $\mathbb{R}[X_{0},X_{1}+X_{2},(1-\sqrt{-1})(X_{1}-\sqrt{-1}X_{2})]$.
-We can show this by a similar argument as above.
-In fact, for all $f(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$, there exists a unique polynomial $\nu(f)(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$ such that
-$f(X_{0},X_{1},X_{2})=\nu(f)(X_{0},X_{1}+X_{2},(1-\sqrt{-1})(X_{1}-\sqrt{-1}X_{2})),$
-and $\beta(f)=f$ if and only if $\nu(f)$ is a real coefficient polynomial.
+By Lemma 6.1, the invariant ring of the homomorphism $\beta\colon\mathbb{C}[X_{0},X_{1},X_{2}]\to\mathbb{C}[X_{0},X_{1},X_{2}]$ that sends $c\in\mathbb{C}$ to $\overline{c}$, $X_{0}$ to $X_{0}$, $X_{1}$ to $X_{2}$ and $X_{2}$ to $X_{1}$ is $\mathbb{R}[X_{0},X_{1}+X_{2},(1-\sqrt{-1})(X_{1}-\sqrt{-1}X_{2})]$.
 Therefore, by setting
 $Y_{0}:=X_{0},\ {Y^{\prime}}_{1}:=X_{1}+X_{2},\ {Y^{\prime}}_{2}:=(1-\sqrt{-1})(X_{1}-\sqrt{-1}X_{2}),$
-and also, setting
+and also, by setting
 $Y_{1}:=\frac{{Y^{\prime}}_{1}}{2},\ Y_{2}:=\frac{{Y^{\prime}}_{2}-{Y^{\prime}}_{1}}{2},$
 we see that the corresponding real form is
-$(Y_{0}^{d}+(Y_{1}+\sqrt{-1})^{d}+(Y_{1}-\sqrt{-1})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{2}.$
-In the surface case, by using the same argument as above, we get the following result:
+$(Y_{0}^{d}+(Y_{1}+\sqrt{-1}Y_{2})^{d}+(Y_{1}-\sqrt{-1}Y_{2})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{2}.$
+In the surface case, by using a similar argument as above, we get the following result:
 Theorem 6.3.
-Assume $d\geq 3,\neq 4$. Then, the correspondence between cocycles and real forms of $F^{2}_{d}$ is given by the following table.
+Assume $d\geq 3,\neq 4$. Then, the correspondence between cocycles and real forms of $F^{2}_{d}$ with odd $d$ is given in the following table:
 cocycle real form
-$K^{0,4}_{2}$ (or $H^{4}_{2}$) $(Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
+$H^{4}_{2}$ $(Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
+$H^{2}_{2}$ $(Y_{0}^{d}+Y_{1}^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
+$H^{0}_{4}$ $((Y_{0}+\sqrt{-1}Y_{1})^{d}+(Y_{0}-\sqrt{-1}Y_{1})^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
+Also, the correspondence between cocycles and real forms of $F^{2}_{d}$ with even $d$ is given in the following table:
+cocycle real form
+$K^{0,4}_{2}$ $(Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
 $K^{1,3}_{2}$ $(-Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
 $K^{2,2}_{2}$ $(-Y_{0}^{d}-Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
-$K^{0,2}_{2}$ (or $H^{2}_{2}$) $(Y_{0}^{d}+Y_{1}^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
+$K^{0,2}_{2}$ $(Y_{0}^{d}+Y_{1}^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
 $K^{1,1}_{2}$ $(-Y_{0}^{d}+Y_{1}^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
-$K^{0,0}_{2}$ (or $H^{0}_{4}$) $((Y_{0}+\sqrt{-1}Y_{1})^{d}+(Y_{0}-\sqrt{-1}Y_{1})^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
-In the same way, we can extend it to the general case. In particular, all real forms of $F_{d}^{n}$ are the $n$-dimensional hypersurface of degree $d$.
+$K^{0,0}_{2}$ $((Y_{0}+\sqrt{-1}Y_{1})^{d}+(Y_{0}-\sqrt{-1}Y_{1})^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
+$L_{2}$ $Z_{d}^{2}$
+Proof.
+The real form corresponding to $L_{2}$ is already described above. First, we consider the real form corresponding to $K_{2}^{k,l}$ with $k+l=4$. The real form corresponding to $K_{2}^{0,4}$ is clear.
+By Lemma 6.1, the invariant ring of the $\mathbb{R}$-homomorphism $\alpha_{1}\colon\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]\to\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]$ that sends $X_{0}$ to $\xi_{d}X_{0}$, $X_{i}$ to $X_{i}$ for $i=1,2,3$, and $c\in\mathbb{C}$ to $\overline{c}$ is $\mathbb{R}[\xi_{2d}X_{0},X_{1},X_{2},X_{3}]$.
+Therefore, by setting
+$Y_{0}:=\xi_{2d}X_{0},\ Y_{i}:=X_{i}\ (i=1,2,3),$
+we see that the real form corresponding to $K_{2}^{1,3}$ is
+$\left(-Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0\right)\subset\mathbb{P}_{\mathbb{R}}^{3}.$
+Also, by Lemma 6.1, the invariant ring of the $\mathbb{R}$-homomorphism $\alpha_{2}\colon\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]\to\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]$ that sends $X_{i}$ to $\xi_{d}X_{i}$ for $i=0,1$, $X_{j}$ to $X_{j}$ for $j=2,3$, and $c\in\mathbb{C}$ to $\overline{c}$ is $\mathbb{R}[\xi_{2d}X_{0},\xi_{2d}X_{1},X_{2},X_{3}]$.
+Therefore, by setting
+$Y_{i}:=\xi_{2d}X_{i}\ (i=0,1),\ Y_{j}:=X_{j}\ (j=2,3),$
+we see that the real form corresponding to $K_{2}^{2,2}$ is
+$\left(-Y_{0}^{d}-Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0\right)\subset\mathbb{P}_{\mathbb{R}}^{3}.$
+Next, we consider the real form corresponding to $K_{2}^{k,l}$ with $k+l=2$.
+By Lemma 6.1, the invariant ring of the $\mathbb{R}$-homomorphism $\beta_{1}\colon\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]\to\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]$ that sends $X_{i}$ to $X_{i}$ for $i=0,1$, $X_{2}$ to $X_{3}$, $X_{3}$ to $X_{2}$, and $c\in\mathbb{C}$ to $\overline{c}$ is $\mathbb{R}[X_{0},X_{1},X_{2}+X_{3},\left(1-\sqrt{-1}\right)\left(X_{2}-\sqrt{-1}X_{3}\right)]$.
+Therefore, by setting
+$Y_{i}:=X_{i}\ (i=0,1),\ {Y^{\prime}}_{2}:=X_{2}+X_{3},{Y^{\prime}}_{3}:=\left(1-\sqrt{-1}\right)\left(X_{2}-\sqrt{-1}X_{3}\right)$
+and also, by setting
+$Y_{2}:=\frac{{Y^{\prime}}_{2}}{2},\ Y_{3}:=\frac{{Y^{\prime}}_{3}-{Y^{\prime}}_{2}}{2},$
+we see that the real form corresponding to $K_{2}^{1,3}$ is
+$\left(Y_{0}^{d}+Y_{1}^{d}+\left(Y_{2}+\sqrt{-1}Y_{3}\right)^{d}+\left(Y_{2}-\sqrt{-1}Y_{3}\right)^{d}=0\right)\subset\mathbb{P}_{\mathbb{R}}^{3}.$
+Also, by Lemma 6.1, the invariant ring of the $\mathbb{R}$-homomorphism $\beta_{2}\colon\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]\to\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]$ that sends $X_{0}$ to $\xi_{d}X_{0}$, $X_{1}$ to $X_{1}$, $X_{2}$ to $X_{3}$, $X_{3}$ to $X_{2}$, and $c\in\mathbb{C}$ to $\overline{c}$ is $\mathbb{R}[\xi_{2d}X_{0},X_{1},X_{2}+X_{3},\left(1-\sqrt{-1}\right)\left(X_{2}-\sqrt{-1}X_{3}\right)]$.
+Therefore, by setting
+$Y_{0}:=\xi_{2d}X_{0},\ Y_{1}:=X_{1},\ {Y^{\prime}}_{2}:=X_{2}+X_{3},{Y^{\prime}}_{3}:=\left(1-\sqrt{-1}\right)\left(X_{2}-\sqrt{-1}X_{3}\right)$
+and also, by setting
+$Y_{2}:=\frac{{Y^{\prime}}_{2}}{2},\ Y_{3}:=\frac{{Y^{\prime}}_{3}-{Y^{\prime}}_{2}}{2},$
+we see that the real form corresponding to $K_{2}^{1,3}$ is
+$\left(-Y_{0}^{d}+Y_{1}^{d}+\left(Y_{2}+\sqrt{-1}Y_{3}\right)^{d}+\left(Y_{2}-\sqrt{-1}Y_{3}\right)^{d}=0\right)\subset\mathbb{P}_{\mathbb{R}}^{3}.$
+Finally, we consider the real form corresponding to $K_{2}^{0,0}$.
+By Lemma 6.1, the invariant ring of the $\mathbb{R}$-homomorphism $\gamma\colon\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]\to\mathbb{C}[X_{0},X_{1},X_{2},X_{3}]$ that sends $X_{i}$ to $X_{i+1}$, $X_{i+1}$ to $X_{i}$ for $i=0,2$, and $c\in\mathbb{C}$ to $\overline{c}$ is $\mathbb{R}[X_{0}+X_{1},\left(1-\sqrt{-1}\right)\left(X_{0}-\sqrt{-1}X_{1}\right),X_{2}+X_{3},\left(1-\sqrt{-1}\right)\left(X_{2}-\sqrt{-1}X_{3}\right)]$.
+Therefore, by setting
+${Y^{\prime}}_{i}:=X_{i}+X_{i+1},{Y^{\prime}}_{i+1}:=\left(1-\sqrt{-1}\right)\left(X_{i}-\sqrt{-1}X_{i+1}\right)\ (i=0,2)$
+and also, by setting
+$Y_{i}:=\frac{{Y^{\prime}}_{i}}{2},\ Y_{i+1}:=\frac{{Y^{\prime}}_{i+1}-{Y^{\prime}}_{i}}{2}\ (i=0,2),$
+we see that the corresponding real form is
+$\left(\left(Y_{0}+\sqrt{-1}Y_{1}\right)^{d}+\left(Y_{0}-\sqrt{-1}Y_{1}\right)^{d}+\left(Y_{2}+\sqrt{-1}Y_{3}\right)^{d}+\left(Y_{2}-\sqrt{-1}Y_{3}\right)^{d}=0\right)\subset\mathbb{P}_{\mathbb{R}}^{3}.$
+In the same way, we can extend it to the general case.
+In particular, all real forms of $F_{d}^{n}$ can be described as the $n$-dimensional hypersurface of degree $d$ except for the real form corresponding to $L_{n}$.
 Next, we will give a topological description of a real locus of these real forms.
 Theorem 6.4.
+  1. (1)
+    The real locus of $Z^{2k}_{d}$ with even $d$ is $\emptyset$ for all $k$ and $d$.
   2. (2)
@@
 In each case, let $Z$ be the real locus.
+(1) As the restriction of $\rho^{\prime}$ to $F^{2k}_{d,\mathbb{C}}$ has no fixed real points, the real locus of $Z^{2k}_{d}$ is $\emptyset$.
 (2) is clear.
-(2) Note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $x_{n+1}\neq 0$. Therefore, we can define a continuous map $Z\to S^{n}$ as follows:
-$[x_{0}:\cdots:x_{n}:1]\mapsto\frac{1}{\sqrt{x_{0}^{2}+\cdots+x_{n}^{2}}}(x_{0},\ldots,x_{n}).$
+(3) Note that $x_{0}:\cdots:x_{n+1}]\in Z$ implies $x_{n+1}\neq 0$. Therefore, we can define a continuous map $Z\to S^{n}$ as follows:
+$x_{0}:\cdots:x_{n}:1]\mapsto\frac{1}{\sqrt{x_{0}^{2}+\cdots+x_{n}^{2}}}(x_{0},\ldots,x_{n}).$
 As it is a bijection, it is also a homeomorphism, and thus, $Z\approx S^{n}$.
-(3) Note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{0},\ldots,x_{i})\neq(0,\ldots,0)$, $(x_{i+1},\ldots,x_{n+1})\neq(0,\ldots,0)$.
+(4) Note that $x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{0},\ldots,x_{i})\neq(0,\ldots,0)$, $(x_{i+1},\ldots,x_{n+1})\neq(0,\ldots,0)$.
 Thus, we can define a continuous map $Z\to\mathbb{P}^{i}(\mathbb{R})\times\mathbb{P}^{n-i}(\mathbb{R})$ as follows:
-$[x_{0}:\cdots:x_{n+1}]\mapsto([x_{0}:\cdots:x_{i}],[x_{i+1}:\cdots:x_{n+1}]).$
+$x_{0}:\cdots:x_{n+1}]\mapsto([x_{0}:\cdots:x_{i}],[x_{i+1}:\cdots:x_{n+1}]).$
 As it is a bijection, it is also a homeomorphism, and thus, $Z\approx\mathbb{P}^{i}(\mathbb{R})\times\mathbb{P}^{n-i}(\mathbb{R})$.
-(4) Note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{1},\ldots,x_{n+1})\neq(0,\ldots,0)$. Therefore, we can define a continuous function $Z\to\mathbb{P}^{n}(\mathbb{R})$ as follows:
-$[x_{0}:\cdots:x_{n+1}]\mapsto[x_{1}:\cdots:x_{n+1}].$
+(5) Note that $x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{1},\ldots,x_{n+1})\neq(0,\ldots,0)$. Therefore, we can define a continuous function $Z\to\mathbb{P}^{n}(\mathbb{R})$ as follows:
+$x_{0}:\cdots:x_{n+1}]\mapsto[x_{1}:\cdots:x_{n+1}].$
 As $d$ is odd, this is a bijection, and thus, a homeomorphism. So $Z\approx\mathbb{P}^{n}(\mathbb{R})$.
-(5) First, note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{n},x_{n+1})\neq(0,0)$. Therefore, we can define a continuous map $Z\to\mathbb{P}^{i}(\mathbb{R})$ as follows:
-$[x_{0}:\cdots:x_{n+1}]\mapsto[x_{n}:x_{n+1}].$
+(6) First, note that $x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{n},x_{n+1})\neq(0,0)$. Therefore, we can define a continuous map $Z\to\mathbb{P}^{i}(\mathbb{R})$ as follows:
+$x_{0}:\cdots:x_{n+1}]\mapsto[x_{n}:x_{n+1}].$
 Then, this image has $\frac{d}{2}$ connected components
 $E_{m}:=\left\{[\cos\theta:\sin\theta]\bigg{|}\frac{(4m-1)\pi}{2d}\leq\theta\leq\frac{(4m+1)\pi}{2d}\right\}$
-for $10\leq m\leq\frac{d}{2}-$ 1.
-In fact, $[x_{n}:x_{n+1}]\in f(Z)$ if and only if $(x_{n}+\sqrt{-1}x_{n+1})^{d}+(x_{n}-\sqrt{-1}x_{n+1})^{d}\leq 0$, and
+for $0\leq m\leq\frac{d}{2}-1$.
+In fact, $x_{n}:x_{n+1}]\in f(Z)$ if and only if $(x_{n}+\sqrt{-1}x_{n+1})^{d}+(x_{n}-\sqrt{-1}x_{n+1})^{d}\leq 0$, and
 $(\cos\theta+\sqrt{-1}\sin\theta)^{d}+(\cos\theta-\sqrt{-1}\sin\theta)^{d}=2\cos d\theta.$
 Also, for each $m$, we can define a continuous map $f^{-1}(E_{m})\to S^{n}$ defined as follows:
-$[x_{0}:\cdots:x_{n-1}:\cos\theta:\sin\theta]\mapsto\frac{1}{\sqrt{x_{0}^{2}+\cdots x_{n-1}^{2}+\sin d\theta}}(x_{0},\ldots,x_{n-1},\sin d\theta)$
+$x_{0}:\cdots:x_{n-1}:\cos\theta:\sin\theta]\mapsto\frac{1}{\sqrt{x_{0}^{2}+\cdots x_{n-1}^{2}+\sin d\theta}}(x_{0},\ldots,x_{n-1},\sin d\theta)$
 Note that this is well-defined as $d$ is even. As this is a bijection, it is also a homeomorphism, and thus $Z\approx\underbrace{S^{n}\sqcup\cdots\sqcup S^{n}}_{\frac{d}{2}}$. ∎
@@
 For this reason, it is also a very challenging problem to determine all real forms of $F^{2}_{4}$ or to see at least if there is a real form of $F^{2}_{4}$ other than these four or not.
-Remark 5.6.
-Let $Q^{r,s}$ be a hypersurface defined as follows:
-$Q^{r,s}:=(X_{0}^{2}+\cdots+X_{r-1}^{2}-X_{r}^{2}-\cdots-X_{r+s-1}^{2}=0)\subset\mathbb{P}^{r+s-1}.$
-Then, $Q^{r,s}$ with $r+s=n+2$ is a real form of $F^{n}_{2}$.
-Also, by considering real loci, we see that $Q^{r,s}$ and $Q^{t,u}$ with $r+s=t+u=n+2$ are $\mathbb{R}$-isomorphic if and only if $r=t$ or $r=u$.
-However, there exists a real form of $F^{n}_{2}$ other than these forms.
-For example, by [Rus02], any real forms of $F^{2}_{2}$ are exactly $\mathbb{R}$-isomorphic to $Q^{4,0}$, $Q^{3,1}$, $Q^{2.2}$, or $Q^{3,0}\times Q^{2,1}$.
+Acknowledgements. I thank Professor János Kollár for pointing out some crucial mistakes and also, giving some advice.